-6x-2y=24
4y=-12x-48

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

-6x-2y=24
4y=-12x-48
-6,-2,24
12,4,-48
-6x-2y=24
12x+4y=-48

no solution
They are the same lines

Answer by stanbon(75887)

-6x-2y=24
4y=-12x-48
----
Rearrange::
2y + 6x = -24
4y + 12x = -48
------
Modify::
4y + 12x = -48
4y + 12x = -48
----
Subtract to get:
0 = 0
------
Solution:: All points of the form (x,-3x-12)
===============
Cheers,
Stan H.
====================